Pinnacle Corp. budgeted $258,700 of overhead cost for the current year. Actual overhead costs for the year were $240,730. Pinnacle's plantwide allocation base, machine hours, was budgeted at 54,860 hours. Actual machine hours were 55,670.

A total of 108,360 units was budgeted to be produced and 98,000 units were actually produced. Pinnacle's plantwide factory overhead rate for the current year is:

a. $4.39 per machine hour b. $4.72 per machine hour c. $2.39 per machine hour d. $2.22 per machine hour

-------------------

Selected accounts with a credit amount omitted are as follows

Work in Process
Apr. 1 Balance 6,500 Apr. 30 Goods finished X
     30 Direct materials 55,700
     30 Direct labor 190,500
     30 Factory overhead 57,150
Finished Goods
Apr. 1 Balance 12,000
     30 Goods finished 309,600

What was the balance of Work in Process as of April 30?

a. $250

b. $309,600

c. $12,000

d. $57,150

-----------------

When Job 117 was completed, direct materials totaled $28,160; direct labor, $33,958; and factory overhead, $20,707. A total of 3,313 units were produced at a per-unit cost of

a. $82,825

b. $3,313

c. $25

d. $62,118

-----------------

Panamint Systems Corporation is estimating activity costs associated with producing disk drives, tapes drives, and wire drives. The indirect labor can be traced to four separate activity pools. The budgeted activity cost and activity base data by product are provided below.


Activity Cost

Activity Base
Procurement $344,000 Number of purchase orders
Scheduling 245,500 Number of production orders
Materials handling 429,500 Number of moves
Product development 727,000 Number of engineering changes
Production 1,445,000 Machine hours

Number of
Purchase
Orders

Number of
Production
Orders

Number
of
Moves

Number of
Engineering
Changes


Machine
Hours

Number
of
Units
Disk drives 4,060    390       1,230 11         2,200 2,000
Tape drives 1,500    135       530 5         8,300 4,500
Wire drives 12,400    910       3,500 28         11,600 2,300

The activity rate for the procurement activity cost pool is

a. $65.38 per purchase order

b. $239.72 per purchase order

c. $171.08 per purchase order

d. $19.15 per purchase order

----------------------

Selected accounts with amounts omitted are as follows

Work in Process
Aug. 1 Balance 277,750 Aug. 31 Goods finished 153,780
     31 Direct materials X
       31 Direct labor 35,400
     31 Factory overhead X
Factory Overhead
Aug. 1 – 31 Costs incurred 96,990 Aug. 1 Balance 15,200
       31 Applied
(30% of direct labor cost) X

If the balance of Work in Process on August 31 is $194,620, what was the amount debited to Work in Process for direct materials in August?

a. $353,990

b. $507,770

c. $118,530

d. $24,630
Answer #1

1) Overhead rate = 258700/54860 = 4.72 per MH

So answer is b) $4.72 per machine hour

2) Ending work in process = 6500+55700+190500+57150-309600 = 250

So answer is b) $250

3) Unit cost

Direct material 28160
Direct labor 33958
Overhead 20707
Total Cost 82825
Unit 3313
Unit cost 25

So answer is c) $25

4) Activity rate for procurement = 344000/17960 = 19.15 per purchase order

So answer is d) $19.15 per purchase order

5) Applied overhead = 35400*30% = 10620

Direct material = 153780+194620-10620-35400-277750 = 24630

So answer is d) $24630